OpenStudy (sheaspease):

Explain how to simplify x(2x + 3x2). You may indicate an exponent in your answer with ^. For example, 3x2 as 3x^2.

OpenStudy (codysek98):

Is it supposed to be x(2x+3x^2)? Or is it 3x2, because that would mean it was x(2x+6x) You would just distribute the x to everything inside of the ( ). Do you know how to multiply exponents?

OpenStudy (sheaspease):

its x(2x+3x2) but they are saying the answer may have ^ in it and not very well

OpenStudy (codysek98):

Well, x*x = x^2. When multiplying, you add the exponents together. For example, 3x^2*x^4 = 3x^6

OpenStudy (codysek98):

To distribute, multiply the x by the first term and then multiply x by the second term and add them together.

OpenStudy (codysek98):

Do you still need help with distributing? x(2x+3x2) = \[2x^2+3x^22\]

OpenStudy (codysek98):

It still confuses me on why they would have the "3x2" as the second term..